# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v1

@@ Abstract
We study a new class of MDPs that employs multinomial logit function approximation to ensure valid probability distributions over the state space. Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency. The best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature space dimension, $H$ is the episode length, and $K$ is the number of episodes. While this result attains the same rate in $K$ as the linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. In this work, we first address the computational issue by proposing an online algorithm that achieves the same regret with only $\mathcal{O}(1)$ computation cost. Then, we design two algorithms that leverage local information to enhance statistical efficiency. They not only maintain an $\mathcal{O}(1)$ computation cost per episode but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ respectively. Finally, we establish a lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for reinforcement learning.
@@ 1 Introduction
Reinforcement Learning (RL) with function approximation has achieved great success in many applications with large state and action space, such as games (Mnih et al., 2013), robotics (Kober et al., 2013), large language models (Ouyang et al., 2022), etc. Therefore, establishing the theoretical foundation for RL with function approximation is of great importance. Recently, there have been many efforts devoted to understanding the linear function approximation due to its significance and simplicity, yielding fruitful results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
While these works make important steps toward understanding reinforcement learning with function approximation, there are still some challenges to be solved. In RL with linear function approximation, transitions are presumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregate probability equals $1$. Thus, for certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$. In general, an algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear as noted by Jin et al. (2020). While there are some works exploring generalized linear function approximation (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Du et al., 2021; Foster et al., 2021), they make assumptions over value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous work and ours in terms of the regret bound, computation cost and storage cost. Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost and storage cost per episode indicate the dependence on episode count $k$.
Reference Regret Comp. Stor. MDP model
@@
UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL+ (Theorem 3) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
@@
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit (MNL) function approximation has been proposed by Hwang and Oh (2023) recently. Despite its benefits, introducing non-linear functions raises significant challenges in both computational and statistical efficiency. Specifically, the best-known method of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret. While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is inefficient. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. Thus, a natural question arises:
Can we achieve (almost) the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation for RL?
@@
In this paper, we answer this question affirmatively by making significant advancements in both computational and statistical efficiency for RL with MNL function approximation, almost matching the results of linear function approximation. Our contributions can be summarized as follows:
@@
  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. It matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
@@
  • As the quantity $\kappa$ can be exponentially small, we propose two algorithms UCRL-MNL-LL and UCRL-MNL-LL+, which leverage local information to improve the statistical efficiency. They not only maintain $\mathcal{O}(1)$ computational and storage costs but also achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, respectively. The latter result moves the $\kappa$ dependence to the lower-order term which is independent of $K$ and the higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in both $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
@@
  • We establish the first lower bound for MDPs with MNL function approximation by employing a reduction to the logistic bandits. We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for the infinite action space setting. Though this does not constitute a strict lower bound for the finite actions case studied in this work, it justifies that our results may be optimal in $d$ and $K$.
Table 1 shows comparisons between previous works and ours. From a technical perspective, inspired by recent advances in multinomial logistic bandits (Zhang and Sugiyama, 2023), we estimate the parameter by running an online algorithm instead of performing maximum likelihood estimation over all historical data. Differently, we need to design value functions carefully that utilize local information to address the unique challenges of MDPs, which is highly non-trivial.
Organization.
@@
We first introduce the related work in Section 2 and present the setup in Section 3. Then, we introduce an algorithm with constant computational and storage costs in Section 4. Next, we present two improved algorithms in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space constraints, we defer all proofs to the appendixes.
Notations.
@@
We denote by $[n]$ the set $\{1,\ldots,n\}$ and define $[x]_{[a,b]}=\min(\max(x,a),b)$. For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. The $\widetilde{\mathcal{O}}(\cdot)$ hides all polylogarithmic factors.
@@ 2 Related Work
In this section, we review related works, both in terms of settings and techniques.
RL with Generalized Linear Function Approximation.
@@
There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is modeled using a multinomial logit model. We significantly improve their results by providing computationally and statistically efficient algorithms.
RL with General Function Approximation.
@@
There have also been some works studies reinforcement learning with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
Multinomial Logit (MNL) Contextual Bandits.
@@
As our work shares similar techniques with MNL contextual bandits to some extent, we briefly review works on this topic. The pioneering study by Cheung and Simchi-Levi (2017) introduced an algorithm that achieves a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). The recent work by Périvier and Goyal (2022) has significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting. Most of the above methods require storing all historical data and are computationally inefficient. A concurrent work with ours by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs inspired by recent advancements in multinomial logistic bandits (Zhang and Sugiyama, 2023) and achieved optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K})$ and $\widetilde{\mathcal{O}}(d\sqrt{K})$ under uniform and nonuniform rewards respectively. However, although the underlying models of MNL contextual bandits and MNL mixture MDPs are similar, the challenges they pose differ significantly and the techniques developed for MNL contextual bandits can not be directly applied to the MNL mixture MDPs. Specifically, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize expected revenue, whereas in MNL mixture MDPs, the goal is to maximize cumulative rewards Therefore, while we utilize similar parameter estimation techniques as those in MNL contextual bandits, we need to construct value functions to address the unique challenges of MDPs, which is highly non-trivial and needs a careful design.
@@ 3 Problem Setup
Inhomogeneous, Episodic MDPs.
An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
@@
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$. The Bellman equation for policy $\pi$ and the Bellman optimality equation are given by
@@
Learning Protocol.
In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$. For the sake of simplicity, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known. The interaction proceeds in $K$ episodes. At the beginning of each episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as the difference between the cumulative reward of the optimal policy and the learner’s policy, namely,
$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}).$
Multinomial Logit (MNL) Mixture MDPs.
Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. We introduce the formal definition of MNL mixture MDPs below.
@@
For any $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, we define the “reachable states” as the set of states that can be reached from state $s$ taking action $a$ at stage $h$ within a single transition, i.e., $\mathcal{S}_{h,s,a}\triangleq\{s^{\prime}\in\mathcal{S}\mid\mathbb{P}_{h}(s^{\prime}\>|\>s,a)>0\}$. Also, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and further define $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
@@
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. Next, we introduce a key assumption for this problem as in Hwang and Oh (2023).

### v2

@@ Abstract
We study a new class of MDPs that employs multinomial logit (MNL) function approximation to ensure valid probability distributions over the state space. Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency. The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes. While this result attains the same rate in $K$ as linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode. Moreover, the quantity $\kappa$ can be exponentially small in the worst case, leading to a significant gap for the regret compared to linear function approximation. In this work, we first address the computational and storage issue by proposing an algorithm that achieves the same regret with only $\mathcal{O}(1)$ cost. Then, we design an enhanced algorithm that leverages local information to enhance statistical efficiency. It not only maintains an $\mathcal{O}(1)$ computation and storage cost per episode but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly closing the gap with linear function approximation. Finally, we establish the first lower bound for MNL function approximation, justifying the optimality of our results in $d$ and $K$.
@@ 1 Introduction
Reinforcement Learning (RL) with function approximation has achieved remarkable success in various applications involving large state and action spaces, such as games (Silver et al., 2016), algorithm discovery (Fawzi et al., 2022) and large language models (Ouyang et al., 2022). Therefore, establishing the theoretical foundation for RL with function approximation is of great importance. Recently, there have been many efforts devoted to understanding the linear function approximation, yielding numerous valuable results (Yang and Wang, 2019; Jin et al., 2020; Ayoub et al., 2020).
While these studies make important steps toward understanding RL with function approximation, there are still some challenges to be solved. In linear function approximation, transitions are assumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregation equals $1$. For certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$ as show by (Hwang and Oh, 2023). While there are also some works exploring generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they made assumptions over value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous works and ours in terms of the regret, computation cost and storage cost. Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost and storage cost per episode indicate the dependence on episode count $k$.
Reference Regret Com. Sto. MDP model
@@
UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
@@
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit function approximation has been proposed by Hwang and Oh (2023) recently. Despite its benefits, the introduction of non-linear functions raises significant challenges in both computational and statistical efficiency. Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret. While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is expensive. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. Thus, a natural question arises:
Is MNL function approximation more difficult than linear function approximation for RL?
@@
In this paper, we answer this question by making significant advancements in both computational and statistical efficiency for MDPs with MNL function approximation, nearly closing the gap with linear function approximation. Table 1 presents a comparison between previous studies and our work. Specifically, our contributions are summarized as follows:
@@
  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. This result matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
@@
  • As the quantity $\kappa$ can be exponentially small, we propose an enhanced algorithm UCRL-MNL-LL, which leverages the local information to improve the statistical efficiency. It not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$. The higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
@@
  • We establish the first lower bound for MDPs with MNL function approximation by introducing a reduction to the logistic bandit problem. We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for MDPs with infinite action space. Though this does not constitute a strict lower bound for the finite actions case studied in this work, it suggests that our result may be optimal with respect to $d$ and $K$.  [^1] After the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) improved the lower bound to $\Omega(dH^{3/2}\sqrt{K})$. This confirms that our result is indeed optimal with respect to $d$ and $K$.
From a technical perspective, inspired by recent advances in logistic bandits (Zhang et al., 2016; Faury et al., 2020; Zhang and Sugiyama, 2023), we observe that the negative log-likelihood function is exponentially concave, which motivates us to apply online Newton step to estimate the unknown transition parameter in an online manner. Moreover, we employ the Bernstein-like inequalities and the self-concordant-like property (Bach, 2010) of the log-loss to achieve the better dependence on $\kappa$.
@@
Organization.  We introduce the related work in Section 2 and present the setup in Section 3. Then, we design an algorithm with constant computational and storage costs in Section 4. Next, we present an algorithm with improved regret guarantee in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space limits, we defer all proofs to the appendixes.
@@
Notations.  Denote by $[n]$ the set $\{1,\ldots,n\}$ and use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$. For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. The $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors.
@@ 2 Related Work
In this section, we review related works from both setup and technical perspectives.
@@
RL with Generalized Linear Function Approximation.  There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is characterized using a multinomial logit model. We significantly improve upon their results by providing computationally and statistically more efficient algorithms.
@@
RL with General Function Approximation.  There have also been some works studies RL with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions to ensure valid probability distributions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
@@
Multinomial Logit Bandits.  There are two types of multinomial logit bandits studied in the literature: the single-parameter model, where the parameter is a vector (Cheung and Simchi-Levi, 2017) and multiple-parameter model, where the parameter is a matrix (Amani and Thrampoulidis, 2021). We focus on the single-parameter model, which are more relevant to our setting. The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds in bandits. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting. Most prior methods required storing all historical data and faced computational challenge. To address this issue, the most recent work by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs building on recent advances in multiple-parameter model (Zhang and Sugiyama, 2023). Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{K}+\kappa^{-1})$ under uniform and non-uniform rewards respectively. However, although the underlying models of MNL bandits and MDPs share similarities, the challenges they present differ substantially, and techniques developed for MNL bandits cannot be directly applied to MNL MDPs. For example, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward. Thus, it is necessary to design new algorithms tailored for MDPs to address these unique challenges.
@@ 3 Problem Setup
Inhomogeneous, Episodic MDPs.  An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
@@
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$. The Bellman equation for policy $\pi$ and Bellman optimality equation are given by
@@
Learning Protocol.  In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$. As learning rewards is no harder than transitions, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known. The interaction proceeds in $K$ episodes. At the beginning of episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as
$\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1}),$
which is the difference between the cumulative reward of the optimal policy and the learner’s policy.
Multinomial Logit (MNL) Mixture MDPs.  Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. We introduce the definition of MNL mixture MDPs below.
@@
For any $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, we define the “reachable states” as the set of states that can be reached from state $s$ taking action $a$ at stage $h$ within a single transition, i.e., $\mathcal{S}_{h,s,a}\triangleq\{s^{\prime}\in\mathcal{S}\mid\mathbb{P}_{h}(s^{\prime}\>|\>s,a)>0\}$. Furthermore, we define $S_{h,s,a}\triangleq|\mathcal{S}_{h,s,a}|$ and denote by $U\triangleq\max_{(h,s,a)}S_{h,s,a}$ the maximum number of reachable states.
Remark 1.
There are many cases that even when the state space is very large, the maximum number of reachable states $U$ is small. This phenomenon is common in situations where the next state is close to the current state. An illustrative example is the RiverSwim problem (Strehl and Littman, 2004).
@@
Remark 2.
While the work of Hwang and Oh (2023) focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$), we address the more general inhomogeneous setting, allowing transitions to vary across different stages.
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. We introduce the key problem-dependent quantities for this problem firstly introduced by Hwang and Oh (2023).